A random sample of 15 items is taken, producing a sample mean of 2.364 with a sample variance of .81. Assume x is normally distributed and construct a 90% confidence interval for the population mean. Appendix A Statistical Tables (Round the answers to 3 decimal places.)

1.672 ≤ μ ≤ 3.056 (wrong)
Answer #1

Solution :

Given that,

t /2,df = 1.761

Margin of error = E = t/2,df * (s /n)

= 1.761 * (0.9 / 15)

Margin of error = E = 0.409

The 90% confidence interval estimate of the population mean is,

- E + E

2.364 - 0.409 2.364 + 0.409

1.955 2.773